OpenStudy (anonymous):

Describe the roots of the equation shown below. x^2-2X+3=0 Two complex roots Two real, irrational roots Two real, rational roots One real, double root

OpenStudy (ranga):

Find the discriminant first. Discriminant = \[b ^{2} - 4ac\] Plug a = 1, b = -2, c = 3 in the above formula Can you tell me what you get for the Discriminant?

OpenStudy (anonymous):

8

OpenStudy (ranga):

Check the sign. That is important.

OpenStudy (ranga):

Discriminant = b^2 - 4ac = (-2)^2 - 4(1)(3) = 4 - 12 = -8 If the discriminant is negative, there will be two complex roots. If the discriminant is zero, there will be two equal real roots. If the discriminant is positive, there will be two real roots.

OpenStudy (anonymous):

-8 sorry

OpenStudy (ranga):

Here the discriminant is negative. Therefore there will be ?

OpenStudy (ranga):

Two complex roots.